Science Report – Nov 6th

Crew 214 Science Report 6 November 2019
SOL: 10
Crew 214 Ð Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The seventh scientific EVA for the mission saw us return to Tank Wash #2 to collect 23 samples from the micrometeorite grid. Four of these had been compromised and it remains to be seen how this will affect the study. 64% of the plots have now been sampled.

With the disappointing failure of the regolith heater prior to arrival at MDRS, a modified plan to demonstrate water extraction was devised. Instead of an in situ test of the device, the gypsum and gypsiferous soil samples collected yesterday would undergo the process in the Science Lab. After documenting the samples, 6 small subsamples of ground gypsum were heated. Within minutes of each subsample being heated, water vapour was observed and video recorded. Further testing of the gypsiferous soil will be attempted tomorrow.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from yesterday’s EVA were processed and experimental procedures commenced.

Closed Loop Food Waste Study: The day’s waste for Sol 10 is: 51 grams; consisting of 175 calories, 4g fat, 2g saturated fat, 328mg sodium, 26g carbohydrates, 3g fibre, 8g sugar, 10g protein.

EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.

EVA Suit Maintenance: Nominal

Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

Copyright © The Mars Society. All rights reserved. | Main Site